OpenStudy (anonymous):

Help please to Solve the equation. If it is an identity (all real numbers) or a contradiction (no solution), so state. 4x – x – 1 = 1 + 3x

OpenStudy (anonymous):

well, what happens when you try to get all the x's on one side?

OpenStudy (anonymous):

move all the x's to the left, and tell me what happens

OpenStudy (anonymous):

ok. brb

OpenStudy (anonymous):

4x – x – 1 = 1 + 3x 4x –x -1 = 1+ 3x 4x- x- 1 -3x = 1 + 3x – 3x -1 1x-x = 0 1x-x/1 = 0/1 x = 0

OpenStudy (anonymous):

that's incorrect

OpenStudy (anonymous):

what is 4x -x?

OpenStudy (anonymous):

4+1

OpenStudy (anonymous):

4x-x = 3x

OpenStudy (anonymous):

it's combining like terms

OpenStudy (anonymous):

we have 4x -x -1 = 1 + 3x

OpenStudy (anonymous):

simplifying the left, we have 3x -1 the right is already simplified

OpenStudy (anonymous):

when we solve equations, our goal is to get the variable, in this case x, on one side so we can solve for it

OpenStudy (anonymous):

we do this by subtracting from both sides, but if we do that here, all the x's disappear entirely

OpenStudy (anonymous):

i know what I'm suppose to be doing but i mostly have problems with identify the signs and putting them in there right order. I subtract when I'm suppose to add the signs

OpenStudy (anonymous):

so you are confusing addition with subtraction?

OpenStudy (anonymous):

at any rate, if all the x's disappear, you are left with 0 = 2

OpenStudy (anonymous):

which is a contradiction

OpenStudy (anonymous):

o ok. So I should do the problem like this: 4x – x – 1 = 1 + 3x 4x- x - 1 -3x = 1+ 3x- 3x 3x /3 = 0/3 x = 0

OpenStudy (anonymous):

Can you please show me on the twiddla board please. I learn better by watching the person do the problem. If you don't mind

OpenStudy (anonymous):

4x – x – 1 = 1 + 3x 4x- x - 1 -3x = 1+ 3x- 3x combine like terms 0 = 1+1 0 = 2

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

twiddle board not necessary.you got it

OpenStudy (anonymous):

the answer is an contradiction because zero can not equal two

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

Thank you so much. Can you please be my tutor? I'm in desperate need of better understanding the basic concepts of the symbols and signs more profoundly.

OpenStudy (anonymous):

I forget about the x turning into 1 that's why I always get it wrong. I have to right that down on a post it next to the computer so I can remember.

OpenStudy (anonymous):

i'm around on the site pretty frequently

OpenStudy (anonymous):

Thank you so much for all your help. I know I really look dumb not remembering this stuff but it's heard trying to focus on math and having small children around and it didn't help that I transferred in this class late. So now I'm trying to play catch up.

OpenStudy (anonymous):

you're not dumb. math requires a lot of time and that's hard to do when you have a life outside of school, but you will only be successful with it if you make the time for it..good luck :)

OpenStudy (anonymous):

Thanks and I will do my to make sure that happens.